# Recommender Systems for Good (RS4Good): Survey of Use Cases and a Call to Action for Research that Matters

### v1

@@ 3. RS4Good: Shifting our Research Focus
Our brief overview in Section 2 demonstrates that already a multitude of RS4Good use cases have been explored in literature. However, the amount of works that focus on RS4Good-related topics is negligible compared to the extensive body of literature devoted to proposing new algorithms exclusively evaluated offline on the few datasets from the media and e-commerce domains. We recall here that there is nothing wrong with research in these areas, and that such research may also fall into the scope of RS4Good. E-commerce recommendations can for example be designed to promote more sustainable products, and a music recommender system’s design may target at increasing the exposure of artists from minority groups. Thus, we iterate that whether a recommender systems serves a ‘for good’ purpose or not rather depends on the design goals of the system than on the application area. However, not many works in the literature address such ‘for good’ purposes in this way.
@@
One part of the answer to the question above may be as simply as this: that it is much easier to do algorithm-oriented research based on offline experiments than to do RS4Good research. Typical RS4Good research can have certain ingredients that may be considered challenging. Usually, one needs to understand the idiosyncrasies of the particular use case, commonly leading to a need for interdisciplinary research. Moreover, the evaluation of RS4Good recommender systems can be demanding, as typically a human-centric research approach is required. Moreover, the aspired positive societal impacts can many times only be assessed in a longitudinal perspective. Research based on offline experiments does not face these challenges. Researchers can, for instance, try out various variants of a newly developed algorithm in a largely automated way without the need to involve humans in experiments. However, the predominant use of offline experimentation in combination problematic offline evaluation practices may have led to a certain level of stagnation in our field.
@@ 3.2.1. Methodological Challenges
    Interdisciplinary: Many of the discussed RS4Good use cases require a deep understanding of the particularities of a given domain. Research in the healthcare domain may be considered a prime example here, which should not rely on a layman’s understanding of real-world issues of a computer scientist. As a result, research in various RS4Good domains requires either interdisciplinary collaboration with experts from other domains or at least an in-depth understanding of the idiosyncrasies the domain.
@@
    Human-in-the-Loop and Impact-Oriented Evaluation: The true value of RS4Good research for different stakeholders in many cases cannot be assessed through abstract computational measures like precision and recall. We may consider a recommendation-based health intervention or the evaluation of a novel recommendation approach in the e-learning domain as examples. In many of these cases, studies require humans in the loop in the evaluation process and measures (e.g., degree of health improvement or learning success) that measure the impact of the interventions or recommendations on users.
@@
Certainly, not all RS4Good use cases may necessarily face all of these methodological challenges in parallel. There may always be certain types of research without involving humans, for instance, works that are based on simulation approaches like the study in the tourism domain by Piliponyte et al. (Piliponyte et al., 2023) mentioned above. [^10] Conversely, research works that are based on user studies or longitudinal developments, e.g., simulations in the e-commerce domain (Zhang et al., 2020; Ghanem et al., 2022), are not necessarily “for good”. In general, however, research in RS4Good domains might often be much slower and effortful in terms of the required research design and the execution of experiments than traditional offline evaluations. Moreover, research on RS4Good use cases, due to the need to consider domain specifics, may always face the criticism of focusing on a too narrow problem setting, and the generalizability of the findings may be questioned. As a result RS4Good research may suffer from a basic dilemma of recommender systems research. On the one hand, researchers in computer science generally strive to develop generalizable solutions, e.g., algorithms, that work for a variety of use cases. On the other hand, there is a concrete danger that we overgeneralize and make conclusions regarding the effectiveness of a new algorithm in practice based on a limited set of experiments and using a specific set of datasets and abstract, domain-agnostic computational measures.
@@ 3.2.2. Research Culture Barriers
The discussed methodological challenges in most cases will probably imply a slower pace of research and publication. Studies involving human subjects often take significant amounts of time, starting with the time needed to recruit suitable participants. Furthermore, working in interdisciplinary teams, getting approval by ethics boards, understanding domain specifics and collaborating with different institutions (e.g., hospitals and doctors) can be time-consuming and effortful as well. Finally, longitudinal studies by definition require a longer observation horizon. Such a slower pace of publication however seems largely incompatible with the fast-paced publication culture in computer science.
For the area of machine learning, Turing-award laureate Yoshua Bengio therefore suggests to entirely re-think the publication process, which nowadays is almost focusing on conferences (Bengio, 2020). This publication process is highly competitive, fast-paced, and driven by deadlines. As a result, Bengio speculates that for many papers that are submitted to conferences and which are eventually published, there might not have been enough time to thoroughly check them for errors. Furthermore, such papers often appear incremental and may lack depth. Ultimately, he advocates for a ‘slow science’ [^11] https://en.wikipedia.org/wiki/Slow_science approach, and envisions a model where research works are rather submitted to fast-turnaround journals, allowing for multiple iterations for improvements, and conference program committees then select already accepted works for presentation and discussion at conferences. This way, the role of conferences can be change into a place for connecting with other people and for highlighting and discussing the “best and most important ideas” within the scientific community.
@@ 4. Summary and Next Steps
With this opinionated essay, we would like to encourage scholars in the field of recommender systems to more often address problems of societal relevance that matter in practice. Our brief survey of examples of existing RS4Good research demonstrates that there are plenty of opportunities and open challenges in various areas from healthcare to education and individual and societal development. We however find that successful RS4Good research requires a fundamental methodological shift in how we do research, both in terms of our methodology and our publication models.
@@
To foster a research culture that values societal impact, conference organizers could, for instance, consider introducing awards specifically for contributions to societal good. While traditional recognitions, such as Best Paper or Test of Time awards, celebrate technological rigor and lasting relevance in the community, awards for societal impact could highlight research that achieves meaningful and positive change. Such recognitions could play an important role in inspiring the community to pursue impactful RS4Good problems, aligning the achievements of the community with societal needs.
@@
We are aware that such changes may take their time. In our view, raising awareness in the community of the issues outlined in this essay represents an important first step. Furthermore, the organization of topical workshops, seminars, conference tracks and journal special issues may help to increase the attention and refocus future efforts in recommender systems research. Generally, we are confident that the community will soon consider personalized ranking algorithms to be a commodity—also due to the increasing capabilities of Large Language Models—and consider the ranking problem to be solved to the extent that a sufficient number of domain-independent well-performing algorithmic solutions are already available by now. Such a development would then allow us as a community to focus on RS4Good research that has the goal to contribute to a positive development of society.

### v2

@@ 3. RS4Good: Shifting our Research Focus
Our brief overview in Section 2 demonstrates that already a multitude of RS4Good use cases have been explored in literature. However, the amount of works that focus on RS4Good-related topics is negligible compared to the extensive body of literature devoted to proposing new algorithms exclusively evaluated offline on the few datasets from the media and e-commerce domains. [^11] This focus on algorithms and accuracy metrics seems particularly pronounced for central recommender systems publication outlets such as the ACM SIGIR, WWW, KDD and RecSys conference series, as well as for general AI conferences such as IJCAI or AAAI. Research on RS4Good use cases, in contrast, does not frequently find its place at these venues. We re-emphasize here that there is nothing wrong with research in these areas, and that such research may also fall into the scope of RS4Good. E-commerce recommendations can, for instance, be designed to promote more sustainable products, and a music recommender system’s design may target at increasing the exposure of artists from minority groups. Thus, we iterate whether a recommender systems serves a ‘for good’ purpose or not rather depends on the design goals of the system than on the application area. However, not many works in the literature address such ‘for good’ purposes in this way.
@@
One part of the answer to the question above may be as simply as this: that it is much easier to do algorithm-oriented research based on offline experiments than to do RS4Good research. Typical RS4Good research can have certain ingredients that may be considered challenging. Usually, one needs to understand the idiosyncrasies of the particular use case, commonly leading to a need for interdisciplinary research (Keestra et al., 2016). Moreover, the evaluation of RS4Good recommender systems can be demanding, as typically a human-centric research approach is required. Moreover, the aspired positive societal impacts can many times only be assessed in a longitudinal perspective. Research based on offline experiments does not face these challenges. Researchers can, for instance, try out various variants of a newly developed algorithm in a largely automated way without the need to involve humans in experiments. However, the predominant use of offline experimentation in combination problematic offline evaluation practices may have led to a certain level of stagnation in our field. To illustrate the predominance of offline experimental designs, we analyzed all papers on recommender systems that were presented at the top-level ACM SIGIR conference between 2022 and 2024. Among the 65 relevant papers, we found no paper that was based on a user study. Furthermore, we could not identify any work that was based on interdisciplinary research. Previous surveys on certain subareas of recommender systems, e.g., (Klimashevskaia et al., 2024; Jannach and Zanker, 2024; Deldjoo et al., 2023), underline that user-centric research is largely underrepresented.
@@ 3.2.1. Methodological Challenges
    Interdisciplinary: Many of the discussed RS4Good use cases require a deep understanding of the particularities of a given domain. Research in the healthcare domain may be considered a prime example here. A computer science with only layman’s understanding of medicine should not be attempting to develop healthcare solutions without the guidance of experts. As a result, research in various RS4Good domains requires either interdisciplinary collaboration with experts from other domains or at least an in-depth understanding of the idiosyncrasies of the domain.
@@
    Human-in-the-Loop and Impact-Oriented Evaluation: The true value of RS4Good research for different stakeholders in many cases cannot be assessed through abstract computational measures like precision and recall. We may consider a recommendation-based health intervention or the evaluation of a novel recommendation approach in the e-learning domain as examples. In many of these cases, studies require humans in the loop in the evaluation process and appropriate measures (e.g., degree of health improvement or learning success) that help us assess the impact of the interventions or recommendations on users.
@@
An orthogonal problem that seems prevalent in various forms of RS4Good research is that of understanding, defining, and properly supporting the goals that should be achieved through a recommender system, both on an individual and a societal level. To optimally support individual goals like self-actualization described above, a participatory approach seems advisable, where users are asked to explicitly state their goals, as done in (Liang et al., 2023) for the case of a podcast recommender system. On a more general level, Ekstrand and Willemsen (2016) argue that there are many cases where “behaviorism is not enough.” While observed user behavior is valuable for recommendation, the authors advocate for increased user involvement and participation in the entire process from algorithm design to evaluation. Besides the problem of capturing the goals, there might also be trade-off situations to be balanced in goal-setting. Remembering the example from the educational domain described above, it might be tempting to recommend easy courses to students that maximize their short-term success in exams. In the long run, however, a more rounded profile including also challenging materials might be more beneficial both for the student and the society. On the individual level, theory-backed mechanisms for adequate goal setting can be applied to address some of these problems. Several works in the health and personal wellbeing domain for example use the Rasch model and alternative approaches to derive goals and make recommendations that consider both the user’s abilities and the difficulty of certain behavioral actions (Schäfer and Willemsen, 2019; Radha et al., 2016; Torkamaan and Ziegler, 2022).
Certainly, not all RS4Good use cases may necessarily face all of these challenges in parallel. For example, there may always be certain types of research without involving humans, for instance, works that are based on simulation approaches like the study in the tourism domain by Piliponyte et al. (Piliponyte et al., 2023) mentioned above. In general, however, research in RS4Good domains might often be much slower and effortful in terms of the required research design and the execution of experiments than traditional offline evaluations. Moreover, research on RS4Good use cases, due to the need to consider domain specifics, may always face the criticism of focusing on a too narrow problem setting, and the generalizability of the findings may be questioned. As a result RS4Good research may suffer from a basic dilemma of recommender systems research. On the one hand, researchers in computer science generally strive to develop generalizable solutions, e.g., algorithms, that work for a variety of use cases. On the other hand, there is a concrete danger that we overgeneralize and make conclusions regarding the effectiveness of a new algorithm in practice based on a limited set of experiments and using a specific set of datasets and abstract, domain-agnostic computational measures.
@@ 3.2.2. Research Culture Barriers
The discussed methodological challenges in most cases will probably imply a slower pace of research and publication. Studies involving human subjects often take significant amounts of time, starting with the time needed to recruit suitable participants. Furthermore, working in interdisciplinary teams, getting approval by ethics boards, understanding domain specifics and collaborating with different institutions (e.g., hospitals and doctors) can be time-consuming and effortful as well. Finally, longitudinal studies like (Liang and Willemsen, 2022; Porcaro et al., 2024) by definition require a longer observation horizon. Such a slower pace of publication however seems largely incompatible with the fast-paced publication culture in computer science.
For the area of machine learning, Turing-award laureate Yoshua Bengio therefore suggests to entirely re-think the publication process, which nowadays is almost focusing on conferences (Bengio, 2020). This publication process is highly competitive, fast-paced, and driven by deadlines. As a result, Bengio speculates that for many papers that are submitted to conferences and which are eventually published, there might not have been enough time to thoroughly check them for errors. Furthermore, such papers often appear incremental and may lack depth. Ultimately, he advocates for a ‘slow science’ [^14] https://en.wikipedia.org/wiki/Slow_science approach, and envisions a model where research works are rather submitted to fast-turnaround journals, allowing for multiple iterations for improvements, and conference program committees then select already accepted works for presentation and discussion at conferences. This way, the role of conferences can change into a place for connecting with other people and for highlighting and discussing the “best and most important ideas” within the scientific community.
@@ 4. Summary and Next Steps
With this opinionated essay, we would like to encourage scholars in the field of recommender systems to more often address problems of societal relevance that matter in practice. Our brief survey of examples of existing RS4Good research demonstrates that there are plenty of opportunities and open challenges in various areas from healthcare to education and individual and societal development. We however find that successful RS4Good research requires a major methodological shift in how we do research, both in terms of our methodology and our publication models.
In terms of the research methodology, more awareness and advanced competencies are required with respect to successfully conducting interdisciplinary research. Various general frameworks for interdisciplinary research have been put forward over the last decades, and several surveys were conducted to understand certain facets, like drivers and barriers, of interdisciplinary collaboration (Rossini and Porter, 1979; Keestra et al., 2016; Tobi and Kampen, 2017; MacRae and Matuszewski, 2024; Siedlok and Hibbert, 2013). RS4Good research should build upon these insights from different fields. Importantly, interdisciplinary research not only means that researchers of different disciplines work together on a problem, but that the work is also based on a shared conceptual model that integrates theoretical frameworks from these disciplines, relies on methodologies that are not limited to one field, and requires the skills of the involved disciplines multiple times throughout the research process (Aboelela et al., 2007). To further strengthen competencies in such collaborative research endeavours, a stronger emphasis on interdisciplinary research skills in computer science education is needed, as advocated for example in (Dabu, 2017). [^15] Notably, the potential benefits of considering and organizing computer sciences as an interdisciplinary study were discussed decades ago in (Agresti, 1976).
Also from a methodological viewpoint, considerable knowledge has been accumulated in different fields regarding appropriate research practices for conducting longitudinal studies. Notable examples can in particular be found in fields such as psychology and medicine (Schaie and Hertzog, 1982; Khoo et al., 2006). In (Lynn, 2009), methods for longitudinal surveys are covered in depth, touching upon a variety of topics like sample design, statistical methods, limitations of longitudinal surveys or ethical questions. Statistical methods for analyzing data that comprise repeated measures for the same individuals over time are covered in depth in (Fitzmaurice and Ravichandran, 2008) and in (Locascio and Atri, 2011), both works from the medical domain. Overall, various methodologies for conducting longitudinal research have been developed in different fields, which can be applied or adapted for RS4Good research.
@@
To foster a research culture and publication models that value societal impact, conference organizers could, for instance, consider introducing awards specifically for contributions to societal good. While traditional recognitions, such as Best Paper or Test of Time awards, celebrate technological rigor and lasting relevance in the community, awards for societal impact could highlight research that achieves meaningful and positive change. Such recognitions could play an important role in inspiring the community to pursue impactful RS4Good problems, aligning the achievements of the community with societal needs.
We are aware that such changes may take their time. In our view, raising awareness on the issues outlined in this essay represents an important first step. In this context, the organization of topical workshops, seminars, conference tracks and journal special issues may help to increase the attention on RS4Good topics and refocus future efforts in recommender systems research. Importantly, such venues should be designed to encourage collaboration across disciplines and provide a dedicated space in particular for human-centric works, case studies and longitudinal research.
@@
Generally, we are confident that the community will soon consider personalized ranking algorithms to be a commodity—also due to the increasing capabilities of Large Language Models—and consider the ranking problem to be solved to the extent that a sufficient number of domain-independent well-performing algorithmic solutions are already available by now. Such a development would then allow us as a community to focus on RS4Good research that has the goal to contribute to a positive development of society.